OpenStudy (anonymous):

Solve using substitution: 4x − y − 4 = 0 7y = 29 + 9x Answer Infinitely many solutions No solution (3, 8) (5, 16)

hero (hero):

Hint: y = 4x - 4

OpenStudy (anonymous):



OpenStudy (anonymous):

thanks so what abt Solve using substitution: -x + 2y = -7 10y − 5x = 21 Answer Infinitely many solutions No solution (-1, -4) (8, 1/2)

OpenStudy (anonymous):

i dnt kn how to do it

OpenStudy (anonymous):

So substitution is when you solve one of the equations for a variable and put it into the other equation

OpenStudy (anonymous):

ok i gt da part

OpenStudy (anonymous):

For this one you could solve -x + 2y = -7 for x. You would add x to both sides and the 7 to both sides to get 2y +7 = x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then put x= 2y + 7 into 10y - 5x = 21. that would look like 10y - 5(2y +7) = 21

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Then simplify. 10y - 5(2y + 7) = 21 10y - 10y - 35 = 21 -35 = 21 NOT TRUE This means there are no solutions. And if you graph then on a calculator you will see the lines are parallel

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

did that make more sense?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

good:)